# Engineering and Computer Science, DGIST, South Korea Bridging Geometric and Semantic Foundation Models

### v1

@@ 5.2 Experimental Setup
MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun] ResNeXt-101 0.236 0.630 0.111 0.885 0.184 0.752 0.332 0.715
DPT [Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun] ViT-Large 0.100 0.901 0.098 0.903 0.078 0.946 0.182 0.758
Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] Stable Diffusion 0.099 0.916 0.055 0.964 0.065 0.960 0.308 0.773
GenPercept [Xu et al.(2024)Xu, Ge, Liu, Fan, Xie, Zhao, Chen, and Shen] Stable Diffusion 0.080 0.934 0.058 0.969 0.096 0.959 0.226 0.741
DepthAnything-V1-Base [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] ViT-Base 0.080 0.941 0.045 0.979 0.055 0.978 0.287 0.754
@@
DepthAnything-V1-Large [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] ViT-Large 0.080 0.947 0.042 0.981 0.057 0.982 0.260 0.758
@@
DepthAnything-V2-Base [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] ViT-Base 0.084 0.936 0.046 0.977 0.062 0.980 0.281 0.75
@@
DepthAnything-V2-Large [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] ViT-Large 0.080 0.944 0.043 0.979 0.053 0.983 0.265 0.754
@@
Table 1: Quantitative comparison of zero-shot relative depth estimation. We compare with the state-of-the-art models that estimate affine-invariant depth. Our evaluation protocol aligns with Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler]. (Bold: best, Underlined: second best)
@@
Table 2: Quantitative comparison of zero-shot depth estimation. Performance on DA-2K [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] benchmark. All metrics are accuracy (%). DA indicates DepthAnything. (Bold: best)
@@
Evaluation Details. We perform zero-shot relative depth evaluation on four datasets: KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun], NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus], ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger], and DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]. Since our method predicts affine-invariant inverse depth, we restrict comparison to relative depth estimation models in this evaluation to ensure a fair comparison. Evaluation is performed by aligning the estimated depth $\mathbf{d}^{*}$ with the ground truth $\mathbf{d}$ by computing scale and shift values. These values are then used to derive an aligned depth map $\tilde{\mathbf{d}}~{}(=\text{scale}\times\mathbf{d}^{*}+\text{shift})$, which is adjusted to the same unit as the ground truth. For quantitative assessment, we use the aligned depth and the ground truth, employing metrics such as AbsRel (absolute relative error: $\frac{1}{M}\sum_{i=1}^{M}\left|\tilde{\mathbf{d}}_{i}-\mathbf{d}_{i}\right|/\mathbf{d}_{i}$) and $\delta_{1}$ accuracy (percentage of pixels where $\max\left(\tilde{\mathbf{d}}_{i}/\mathbf{d}_{i},\mathbf{d}_{i}/\tilde{\mathbf{d}}_{i}\right)<1.25$). We further evaluate our method on DA-2K [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao], a high-resolution benchmark comprising diverse scenes, which is well-suited for analyzing fine-grained details and generalization performance. In this benchmark, the accuracy is measured by identifying the closer of two distinct pixels. Given this evaluation protocol based on pairwise depth comparison, we extend our evaluation to include metric depth estimation models, enabling a comprehensive assessment of our method.
5.3 Comparison to State-of-the-art Methods
We evaluate our BriGeS on zero-shot depth estimation through both quantitative and qualitative comparisons. Table 1 provides a comparison of our model with six state-of-the-art relative depth estimation methods including MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun], DPT [Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun], Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler], GenPercept [Xu et al.(2024)Xu, Ge, Liu, Fan, Xie, Zhao, Chen, and Shen], DepthAnything-V1 [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] and V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] across multiple benchmark datasets [Geiger et al.(2012)Geiger, Lenz, and Urtasun, Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus, Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger, Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]. For the DA-2K benchmark [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao], we extend the evaluation to include nine representative methods, covering both relative and metric depth estimation approaches. In addition to six models above, we include UniDepth [Piccinelli et al.(2024)Piccinelli, Yang, Sakaridis, Segu, Li, Van Gool, and Yu], Metric3d-V2 [Hu et al.(2024)Hu, Yin, Zhang, Cai, Long, Chen, Wang, Yu, Shen, and Shen], and DepthPro [Bochkovskii et al.(2025)Bochkovskii, Delaunoy, Germain, Santos, Zhou, Richter, and Koltun], as detailed in Table 2.
Figure 3: Qualitative results on unseen datasets of BriGeS (Ours) and various state-of-the-art methods. BriGeS is highly effective at capturing intricate structures and delicate objects, such as thin wires, tree branches, and nets. Best viewed when zoomed in.
Quantitative Comparison. As shown in Table 1, BriGeS consistently achieves enhanced performance across the majority of datasets. Compared to DepthAnything-V1 and V2, it achieves an average reduction of 7.33% in the AbsRel metric. Notably, the most significant performance improvement is observed in the DIODE dataset [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.], with an AbsRel reduction of 15.33% in the BriGeS + DepthAnything-V1-Base setup, highlighting its remarkable effectiveness. Similarly, in Table 2, BriGeS achieves improved performance on the DA-2K benchmark.
Temperature Scaling Factor 2 2.5 3 3.5 4
Avg. Rank (↓) 2.3 1.3 2 1.5 2.1
Table 3: Ablation study on the scaling factor $\tau$ in our Attention Temperature Scaling. (Bold: best)
[…]
figureQualitative results of ablation study on proposed modules. A reversed color map is applied for better visualization.
When built upon DepthAnything-V2, it achieves the highest performance on the DA-2K, surpassing other state-of-the-art methods, demonstrating that integrating geometric and semantic information significantly enhances depth estimation accuracy.
Qualitative Comparison. Fig. 1, 3 show representative depth estimation results from our method. These collectively indicate that existing models generally fail to capture semantic information for accurate depth estimation. The first row from Fig. 3 showcases the inference results of thin power lines, which are best preserved in our method, demonstrating its capacity to recover delicate structures. Similarly, in the second row, our model captures the intricate and thin structure of tree branches more effectively than DepthAnything-V2, while in the third row, it best preserves the fine-grained details of the fishing net. Although Marigold and GenPercept, which leverage Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], perform well with fine details, they mispredict the sky, as seen in the first and second rows. These results show that our BriGeS effectively fuses geometric and semantic information, accurately delineating object edges at similar depths and preserving fine details in complex structures.
5.4 Ablation Study
We conduct ablation studies on BriGeS using both DepthAnything-V1-Base and Large [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao] to assess the effectiveness and robustness of our proposed methods across different scales.
The quantitative results are presented in Table 4 for the base and large models, showcasing consistent performance improvements. As evident in Table 4-(b), our Bridging Gate effectively compensates for the lack of semantic information in the geometric features across different model sizes. This improvement is consistently observed in both models, demonstrating the versatility and effectiveness of our Bridging Gate in enhancing semantic details irrespective of the model’s scale.
Model size # Component KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun] NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus] ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger] DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]
@@ 5.4 Ablation Study
Table 4: Ablation study on proposed modules. (a) The baseline model DepthAnything-V1 [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao]. (b) The baseline with Bridging Gate. (c) Our method includes the Bridging Gate and the Attention Temperature Scaling. (Bold: best, underlined: second best)
@@
Additionally, Table 4-(c) demonstrates that our Attention Temperature Scaling effectively enhances the model’s overall performance in both model sizes. These results confirm that the highest performance is achieved when all proposed modules are applied, demonstrating the synergistic effect of our approaches.
@@
Table 3 presents the average rank calculated for the quantitative results across KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun], NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus], ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger], and DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.] by varying the temperature scaling factor for BriGeS based on DepthAnything-V1-Large. We empirically set the scaling factor to 2.5, as it yields the best performance. The qualitative results in Fig. 3 further underscore the impact of our methods. Notably, without the scaling module, the tower cranes within the black box were not well predicted, an issue that is effectively resolved with the scaling module. These results demonstrate that the Bridging Gate and the scaling module help distribute attention more effectively. By mitigating excessive focus on central objects, they reduce prediction errors associated with small structures, thereby enhancing overall performance.
6 Conclusion
@@
Supplementary Material
Appendix A More Qualitative Results
For a detailed examination of the qualitative performance across four unseen datasets, please refer to Fig. 4 through Fig. 7. These figures present the results for the KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun], NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus], ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger], DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.], respectively. In this section, we compare our model with DepthAnything-V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]. Our model outperforms particularly in complex scenes, such as those containing intricate structures and thin objects.
Figure 4: Qualitative results on KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun].
Figure 5: Qualitative results on NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus].
Figure 6: Qualitative results on ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger].
Figure 7: Qualitative results on DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.].

### v2

@@ III Experiments
MiDaS ResNeXt-101 0.236 0.630 0.111 0.885 0.184 0.752 0.332 0.715
DPT ViT-Large 0.100 0.901 0.098 0.903 0.078 0.946 0.182 0.758
Marigold Stable Diffusion 0.099 0.916 0.055 0.964 0.065 0.960 0.308 0.773
GenPercept Stable Diffusion 0.080 0.934 0.058 0.969 0.096 0.959 0.226 0.741
DepthAnything-V1-Base ViT-Base 0.080 0.941 0.045 0.979 0.055 0.978 0.287 0.754
@@
DepthAnything-V1-Large ViT-Large 0.080 0.947 0.042 0.981 0.057 0.982 0.260 0.758
@@
DepthAnything-V2-Base ViT-Base 0.084 0.936 0.046 0.977 0.062 0.980 0.281 0.750
@@
DepthAnything-V2-Large ViT-Large 0.080 0.944 0.043 0.979 0.053 0.983 0.265 0.754
@@
TABLE II: Comparison of zero-shot depth estimation on DA-2K [24]. All metrics are accuracy (%). (Bold: best, DA: DepthAnything)
Relative depth estimation model
MiDaS [14] DPT [13] Marigold [8] GenPercept [22] DA-V1-B [23] DA-V1-L [23] DA-V2-B [24] DA-V2-L [24]
83.5 79.1 86.8 93.2 89.4 88.5 97.0 97.1
Metric depth estimation model BriGeS (Ours)
UniDepth [12] Metric3D-V2 [7] DepthPro [1] DA-V1-B DA-V1-L DA-V2-B DA-V2-L
83.7 89.2 94.7 91.4 92.1 97.0 97.6
III-A Implementation Details
Our training methodology aligns closely with the one employed in the DepthAnything [23, 24]. Initially, the ground truth depth value is transformed into a disparity space and then normalized to a range of 0 to 1. To effectively train on datasets with varying depths, we employ an affine-invariant loss, which ignores the effects of scale and shift. When training BriGeS with DepthAnything-V1, only the affine-invariant loss is used, whereas training with DepthAnything-V2 incorporates both the affine-invariant loss and the gradient matching loss from MiDaS [14] at a 1:2 ratio.
The batch size is set to 16, with an initial learning rate of 5e55e-5. We use the AdamW optimizer [11] with a linear schedule. We conduct experiments on 8 RTX TITAN GPUs using labeled images, running for 2 epochs with DepthAnything-V1 and 1 epoch with DepthAnything-V2, respectively. Training our model for a single epoch takes 8 hours for DepthAnything-Base and 18 hours for DepthAnything-Large. During training, all images are resized so that the shorter side is 518 while maintaining the original aspect ratio, and then randomly cropped to a size of $\times518$. In addition, the temperature scaling factor $\tau$ is set to 2.5 during the inference. Our model contains 244M parameters based on the DepthAnything-Base, and 744M based on the DepthAnything-Large, including the SegmentAnything [9] encoder.
III-B Experimental Setup
Training Datasets. For training BriGeS, we employ a carefully curated subset of the extensive datasets featured in DepthAnything [23, 24]. To enhance the model’s generalization capabilities, we select three diverse datasets encompassing both indoor and outdoor environments for each model. Our training data comprises 600K samples from BlendedMVS [25], HRWSI [21], and TartanAir [20] for DepthAnything-V1, and 582K samples from VKITTI [3], Hypersim [15], and TartanAir for DepthAnything-V2, which corresponds to approximately 1% of the training datasets originally used to train the respective DepthAnything.
@@ III-B Experimental Setup
Evaluation Details. We perform zero-shot relative depth evaluation on four datasets: KITTI [5], NYUv2 [18], ETH3D [17], and DIODE [19]. Since our method predicts affine-invariant inverse depth, we restrict comparison to relative depth estimation models in this evaluation to ensure a fair comparison. Evaluation is performed by aligning the estimated depth $\mathbf{d}^{*}$ with the ground truth $\mathbf{d}$ by computing scale and shift values. These values are then used to derive an aligned depth map $\tilde{\mathbf{d}}~(=\text{scale}\times\mathbf{d}^{*}+\text{shift})$, which is adjusted to the same unit as the ground truth. For quantitative assessment, we use the aligned depth and the ground truth, employing metrics such as AbsRel (absolute relative error: $\frac{1}{M}\sum_{i=1}^{M}\left|\tilde{\mathbf{d}}_{i}-\mathbf{d}_{i}\right|/\mathbf{d}_{i}$) and $\delta_{1}$ accuracy (percentage of pixels where $\max\left(\tilde{\mathbf{d}}_{i}/\mathbf{d}_{i},\mathbf{d}_{i}/\tilde{\mathbf{d}}_{i}\right)<1.25$). We further evaluate our method on DA-2K [24], a high-resolution benchmark comprising diverse scenes, which is well-suited for analyzing fine-grained details and generalization performance. In this benchmark, the accuracy is measured by identifying the closer of two distinct pixels. Given this evaluation protocol based on pairwise depth comparison, we extend our evaluation to include metric depth estimation models, enabling a comprehensive assessment of our method.
Figure 3: Qualitative results on unseen datasets of BriGeS (Ours) and various state-of-the-art methods.
TABLE III: Ablation study on proposed modules. (a) The baseline model DepthAnything-V1 [23]. (b) The baseline with Bridging Gate. (c) Our method includes the Bridging Gate and the Attention Temperature Scaling. (Bold: best, underlined: second best)
Model size # Component KITTI [5] NYUv2 [18] ETH3D [17] DIODE [19]
@@
III-C Comparison to State-of-the-art Methods
We evaluate our BriGeS on zero-shot depth estimation through both quantitative and qualitative comparisons. Table I provides a comparison of our model with six state-of-the-art relative depth estimation methods including MiDaS [14], DPT [13], Marigold [8], GenPercept [22], DepthAnything-V1 [23] and V2 [24] across multiple benchmark datasets [5, 18, 17, 19]. For the DA-2K benchmark [24], we extend the evaluation to include nine representative methods, covering both relative and metric depth estimation approaches. In addition to six models above, we include UniDepth [12], Metric3d-V2 [7], and DepthPro [1], as detailed in Table II.
Quantitative Comparison. As shown in Table I, BriGeS consistently achieves enhanced performance across the majority of datasets. Compared to DepthAnything-V1 and V2, it achieves an average reduction of 7.33% in the AbsRel metric. Notably, the most significant performance improvement is observed in the DIODE dataset [19], with an AbsRel reduction of 15.33% in the BriGeS + DepthAnything-V1-Base setup, highlighting its remarkable effectiveness. Similarly, in Table II, BriGeS achieves improved performance on the DA-2K benchmark. When built upon DepthAnything-V2, it achieves the highest performance on the DA-2K, surpassing other state-of-the-art methods, demonstrating that integrating geometric and semantic information significantly enhances depth estimation accuracy.
TABLE IV: Ablation study on the scaling factor $\tau$ in our Attention Temperature Scaling. (Bold: best)
Temperature Scaling Factor 2 2.5 3 3.5 4
Avg. Rank (↓) 2.3 1.3 2 1.5 2.1
Figure 4: Qualitative results of ablation study on proposed modules.
Qualitative Comparison. Fig. 1, 3 show representative depth estimation results from our method. These collectively indicate that existing models generally fail to capture semantic information for accurate depth estimation. The first row from Fig. 3 showcases the inference results of thin power lines, which are best preserved in our method, demonstrating its capacity to recover delicate structures. Similarly, in the second row, our model captures the intricate and thin structure of tree branches more effectively than DepthAnything-V2, while in the third row, it best preserves the fine-grained details of the fishing net. Although Marigold and GenPercept, which leverage Stable Diffusion [16], perform well with fine details, they mispredict the sky, as seen in the first and second rows. These results show that our BriGeS effectively fuses geometric and semantic information, accurately delineating object edges at similar depths and preserving fine details in complex structures.
III-D Ablation Study
@@
We conduct ablation studies on BriGeS using both DepthAnything-V1-Base and Large [23] to assess the effectiveness and robustness of our proposed methods across different scales. The quantitative results are presented in Table III for the base and large models, showcasing consistent performance improvements. As evident in Table III-(b), our Bridging Gate effectively compensates for the lack of semantic information in the geometric features across different model sizes. This improvement is consistently observed in both models, demonstrating the versatility and effectiveness of our Bridging Gate in enhancing semantic details irrespective of the model’s scale. Additionally, Table III-(c) demonstrates that our Attention Temperature Scaling effectively enhances the model’s overall performance in both model sizes. These results confirm that the highest performance is achieved when all proposed modules are applied, demonstrating the synergistic effect of our approaches.
@@
Table IV presents the average rank calculated for the quantitative results across KITTI [5], NYUv2 [18], ETH3D [17], and DIODE [19] by varying the temperature scaling factor for BriGeS based on DepthAnything-V1-Large. We empirically set the scaling factor to 2.5, as it yields the best performance. The qualitative results in Fig. 4 further underscore the impact of our methods. Notably, without the scaling module, the tower cranes within the black box were not well predicted, an issue that is effectively resolved with the scaling module. These results demonstrate that the Bridging Gate and the scaling module help distribute attention more effectively. By mitigating excessive focus on central objects, they reduce prediction errors associated with small structures, thereby enhancing overall performance.
IV Conclusion